OpenStudy (dls):

A constant power is supplied to a rotating disc.Angular velocity(w) of a disc varies with number of rotations (n) made by the disc as :

OpenStudy (dls):

@Mashy !!!

OpenStudy (dls):

@yrelhan4

OpenStudy (anonymous):

\[p= \tau. \theta \]

OpenStudy (dls):

yes thats constant

OpenStudy (dls):

isnt it tau * omega?

OpenStudy (anonymous):

sorry .. its omega!!

OpenStudy (dls):

:|

OpenStudy (anonymous):

wait lemme think logically!

OpenStudy (dls):

(Y)

OpenStudy (anonymous):

wait.. you can't use that :D.. thats only instantaneous power.. or when angular velocity ain't changing.. like for example when friction is present!

OpenStudy (dls):

so?

OpenStudy (anonymous):

so if power is constant.. work done per unit time is constant hence \[\tau \theta \] is a constant per unit time

OpenStudy (dls):

okay!

OpenStudy (anonymous):

wait wait.. this ain't making sense.. lets go work energy theorem work done in lets say t seconds = change in kinetic energy!.. i think we can use that and get the answer!

OpenStudy (dls):

try it :P

OpenStudy (anonymous):

ok pt = change in kinetic energy =\[\frac {1} {2} I(\omega_2^2 - \omega_1^2) = \tau \theta\]

OpenStudy (anonymous):

total number of rotations n = 2 pi / theta

OpenStudy (anonymous):

so is omega proportional to (theta)^ (-1/2) :D

OpenStudy (dls):

We have to tell the relation b/w omega and n

OpenStudy (anonymous):

i mean't n.. not theta :P.. so i guess its wrong

OpenStudy (dls):

obviously wrong :P

OpenStudy (dls):

Options are : omega propotional to n^1/3 omega propotional to n^3/2 omega propotional to n^2/3 omega propotional to n^2

OpenStudy (anonymous):

and n= theta/2pi.. so my earlier answer is power 1/2... but thats not in the options :P damn.. wait.. lets see

OpenStudy (anonymous):

ok wait.. since we need in terms of n n is total number of rotations FROM beginning.. so we need to consider at any time t from the beginning so \[n = \frac{\theta }{2 \pi }=> \theta = 2n \pi \] total work done = Pt = \[\frac{1}{2} I \omega^2 = \tau \theta = 2n \pi \] n is number of rotations or number of rotations per second?

OpenStudy (dls):

i guess so

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

thinking logically i feel this if power is constant.. that means force must be constant.. if a constant force is put.. then usually velocity linearly increases with time and displacement increases exponentially with time so w cannot be proportional to n power something bigger than one.. so rule out option 2 and four :D.. plz tell me m right in ruling out those options?

OpenStudy (dls):

yes :P

OpenStudy (dls):

\[\LARGE \alpha(\omega.\frac{d \omega}{d \theta})\omega=P\]

OpenStudy (dls):

just take a break and tell me :P  what will be velocity at point X??